OpenStudy (anonymous):

Which is the equation of the line, in point-slope form, that has a slope of 12 and passes through the point (–1, 14)?

OpenStudy (texaschic101):

point slope formula : y - y1 = m(x - x1) slope(m) = 12 (-1,14) x1 = -1 and y1 = 14 now sub your info into the formula and you will have the equation in point slope form. You do not have to solve it, just plug your stuff into the formula. Your answer will still have the parenthesis in it.